# thanks: corresponding author

### v3

@@ 1 Introduction
$\bm{\rho}(i_{1}\cdots i_{n},j_{1}\cdots j_{n})=\bm{X}_{1}^{i_{1},j_{1}}\bm{X}_{2}^{i_{2},j_{2}}\cdots\bm{X}_{n}^{i_{n},j_{n}},$ (1)
@@ 1.1 Main contributions
Table 1: Comparison with existing work for estimating $n$-qudits MPOs (with bond dimension $\overline{r}$) to achieve $\epsilon$-closeness in the Frobenius norm. The result for SIC/2-designs depends on $\gamma(\bm{\rho})=K\cdot\max_{k}p_{k}$, i.e., the probability distribution of measuring the state $\bm{\rho}$ with the corresponding POVM.
POVM ##\## POVMs ##\## total state copies
@@
SIC/$2$-designs (Theorem 1)             $1$ $O(\frac{nd^{2}\gamma(\bm{\rho})\overline{r}^{2}\log n}{\epsilon^{2}})$
$t$-designs, $t\geq 3$ (Theorem 2)             $1$ $O(\frac{nd^{2}\overline{r}^{2}\log n}{\epsilon^{2}})$
@@
We begin by studying recovery guarantees for MPOs from empirical quantum measurements (physical measurements containing statistical errors) obtained using a class of IC-PVOMs including the rank-one symmetric IC-POVM (SIC-POVM) [38] and more broadly spherical $t$-design POVMs that are induced by (approximate) spherical $t$-designs [39, 40]. The first contribution of this paper—presented in Section 3—is a new sample complexity bound (refers to the number of state copies) that matches the number of parameters in MPOs to guarantee bounded recovery error for a particular estimator—the solution to a constrained least-squares optimization problem. Specifically, for SIC-POVM and $2$-design POVMs, our result shows that $M=O(nd^{2}\gamma(\bm{\rho})\overline{r}^{2}\log n/\epsilon^{2})$ number of state copies is sufficient to achieve $\epsilon$-accuracy in recovering the ground-truth state $\bm{\rho}$ in the Frobenius norm, where $\gamma(\bm{\rho})=K\cdot\max_{k=1}^{K}p_{k}$ captures the uniformity of the probabilistic distribution of the measurement. The term $\gamma(\bm{\rho})$ can be as small as 1, but in some cases, it may be as large as $d^{n}$; however, for many cases it is only polynomially large in $n$. The dependence of $\gamma(\bm{\rho})$ could be eliminated by using spherical $t$-designs with $t\geq 3$, for which we only need $M=O(nd^{2}\overline{r}^{2}\log n/\epsilon^{2})$ to ensure the sufficient condition. Excluding the logarithmic term, this bound exhibits linear dependence on $n$, $d^{2}$, and $\overline{r}^{2}$, aligning precisely with the degrees of freedom $O(nd^{2}\overline{r}^{2})$ in MPOs and is thus sample efficient. This result improves upon previous work [33] mentioned above using Haar random projective measurement bases that requires $\Omega(nd^{2}\overline{r}^{2}\log n)$ POVMs and $O(n^{3}d^{2}\overline{r}^{2}\log n/\epsilon^{2})$ total number of state copies. We summarize the main results and comparison with our previous result [33] in Table 1. Moreover, physical MPO states are often approximately low-rank, allowing the previous analysis to be extended to the trace distance. Specifically, to ensure a bounded recovery error in terms of trace distance for rank-$s$ MPO states, the required number of state copies remains proportional to the degrees of freedom, $O(nsd^{2}\overline{r}^{2})$, when using (approximate) spherical $t$-designs with $t\geq 2$.
@@
The second contribution of this paper—presented in Section 4—is the development of a projected gradient descent (PGD) with guaranteed convergence analysis for solving the aforementioned constrained least-squares optimization problem. In each iteration, the algorithm performs a gradient update, followed by a projection back to the MPO format with a unit trace. Even without the constraint on the trace, computing the optimal projection onto the set of MPOs is challenging and could be NP-hard [41]. We employ a sequential approach called TT-SVD [27]—developed for computing TT decomposition which is equivalent to MPO—for computing an approximate projection. While TT-SVD is not optimal, it is computationally efficient and comes with a sub-optimal guarantee. Leveraging this guarantee, we show that, with an appropriate initialization, PGD converges linearly to a vicinity of the ground-truth state, with a distance matching the results in Table 1. In addition, we show that spectral initialization provides a valid starting point, ensuring the convergence of PGD. Our numerical experiments in Section 5 suggest that PGD can also efficiently find the underlying state even with a random initialization.
@@
We note that our work is centered on the fundamental question of whether the number of state copies required for QST can be efficient (i.e. proportional to the intrinsic degrees of freedom) for recovering a structured quantum many-body state. We investigate this problem using IC-POVMs, focusing on sample complexity while not accounting for the implementation complexity of the measurement settings. In practice, the SIC-POVM or the spherical $t$-designs with $t\geq 2$ have no known efficient implementation using local quantum circuits. To partially address this practicality concern, we provide numerical evidence that local SIC-POVMs, which can be implemented efficiently and readily on current quantum hardware, can be used to perform QST for typical MPO states with $O(n)$ state copies using the proposed PGD algorithm in Section 5.
1.2 Notation
We use calligraphic letters (e.g., $\mathcal{X}$) to denote tensors, bold capital letters (e.g., $\bm{X}$) to denote matrices, bold lowercase letters (e.g., $\bm{x}$) to denote column vectors, and italic letters (e.g., $x$) to denote scalar quantities. Elements of matrices and tensors are denoted in parentheses. For example, $\mathcal{X}(i_{1},i_{2},i_{3})$ denotes the element in position $(i_{1},i_{2},i_{3})$ of the order-3 tensor $\mathcal{X}$. The calligraphic letter $\mathcal{A}$ is reserved for the linear measurement map. For a positive integer $K$, $[K]$ denotes the set $\{1,\dots,K\}$. The superscripts $(\cdot)^{\top}$ and $(\cdot)^{\dagger}$ denote the transpose and Hermitian transpose, respectively  [^3] As is conventional in the quantum physics literature (but not in information theory and signal processing), we use $(\cdot)^{\dagger}$ to denote the Hermitian transpose.. For two matrices $\bm{A},\bm{B}$ of the same size, $\left\langle\bm{A},\bm{B}\right\rangle=\operatorname{trace}(\bm{A}^{\dagger}\bm{B})$ denotes the inner product between them. $\|\bm{X}\|$ and $\|\bm{X}\|_{F}$ respectively represent the spectral norm and Frobenius norm of $\bm{X}$. $\sigma_{i}(\bm{X})$ is the $i$-th singular value of $\bm{X}$. For a vector $\bm{a}$ of size $N\times 1$, its $l_{n}$-norm is defined as $||\bm{a}||_{n}=(\sum_{m=1}^{N}|a_{m}|^{n})^{\frac{1}{n}}$. For two positive quantities $a,b\in\mathbb{R}$, the inequality $b\lesssim a$ or $b=O(a)$ means $b\leq ca$ for some universal constant $c$; likewise, $b\gtrsim a$ or $b=\Omega(a)$ represents $b\geq ca$ for some universal constant $c$. To simplify notations in the following sections, for an order-$n$ TT format with ranks $(r_{1},\dots,r_{n})$, we define $\overline{r}=\max_{i=1}^{n-1}r_{i}$.
2 Informationally Complete POVM
@@
$\|\mathcal{A}(\bm{\rho})\|_{2}^{2}=\sum_{k=1}^{d^{2n}}\langle\bm{A}_{k},\bm{\rho}\rangle^{2}=\frac{\|\bm{\rho}\|_{F}^{2}+(\operatorname{trace}(\bm{\rho}))^{2}}{d^{n}(d^{n}+1)}.$ (3)
The proof is provided in Appendix A. By using tools from compressive sensing literature [57, 58, 59, 60, 61], the previous work [33] provides a similar stable embedding result from Haar random projective measurement bases, but only for MPOs. In contrast, with POVM being informationally complete, (3) holds for any Hermitian matrices without relying on any randomness in the measurement operator. This allows us to develop tight sample complexity bound in Section 3 and iterative algorithms with guaranteed convergence in Section 4. As an immediate consequence of Lemma 1, for any two density matrices $\bm{\rho}_{1}$ and $\bm{\rho}_{2}$, the difference $\bm{\rho}_{1}-\bm{\rho}_{2}$ remains Hermitian, and hence
@@
(Spherical $t$-designs [62, 15]). A finite set $\{\bm{w}_{k}\}_{k=1}^{K}\subset\mathbb{C}^{d^{n}}$ of normalized vectors is called a spherical quantum $t$-design if 444 $K\geq C_{d^{n}+\lfloor t/2\rfloor-1}^{\lfloor t/2\rfloor}\cdot C_{d^{n}+\lceil t/2\rceil-1}^{\lceil t/2\rceil}$ is necessary to form a spherical $t$-design[52, 63].
@@
[…]
@@
[…]
@@
$(1-\delta)\frac{d^{n}(\|\bm{\rho}\|_{F}^{2}+(\operatorname{trace}({\bm{\rho}}))^{2})}{K(d^{n}+1)}\leq\|\mathcal{A}(\bm{\rho})\|_{2}^{2}$
$\leq(1+\delta)\frac{d^{n}(\|\bm{\rho}\|_{F}^{2}+(\operatorname{trace}({\bm{\rho}}))^{2})}{K(d^{n}+1)}.$ (7)
The proof is given in Appendix B. Similar to (4), it follows from Lemma 2 that for any two density matrices $\bm{\rho}_{1}$ and $\bm{\rho}_{2}$, $\mathcal{A}(\bm{\rho}_{1}-\bm{\rho}_{2})$ satisfies
@@
3 Stable Recovery of Matrix Product Operators
@@
$\widehat{\bm{\rho}}=\operatorname*{\text{arg\leavevmode\nobreak\ min}}_{\bm{\rho}\in\mathbb{X}_{\bm{r}}}\|\mathcal{A}(\bm{\rho})-{\widehat{\bm{p}}}\|_{2}^{2},$ (10)
@@
00 0 $\leq$ $\|\mathcal{A}(\bm{\rho}^{\star})-\widehat{\bm{p}}\|_{2}^{2}-\|\mathcal{A}(\widehat{\bm{\rho}})-\widehat{\bm{p}}\|_{2}^{2}$ (12)
@@
$\operatorname{\mathbb{P}}\left(\sum_{k=1}^{K}a_{k}(\frac{f_{k}}{M}-p_{k})>t\right)$
$\leq e^{-\frac{Mt}{4a_{\max}}\min\bigg{\{}1,\frac{a_{\max}t}{4\sum_{k=1}^{K}a_{k}^{2}p_{k}}\bigg{\}}}+e^{-\frac{Mt^{2}}{8\sum_{k=1}^{K}a_{k}^{2}p_{k}}},$
@@
Case $(i)$: $2$-designs.
There are two potential technical difficulties in directly applying Lemma 3 to bound the term $\langle\bm{\eta},\mathcal{A}(\widehat{\bm{\rho}}-\bm{\rho}^{\star})\rangle$ by plugging in $a_{k}=\langle\bm{A}_{k},\widehat{\bm{\rho}}-\bm{\rho}^{\star}\rangle$: (1) $a_{k}$ in not independent to the random variables $\{f_{k}\}$ since $\widehat{\bm{\rho}}$ depends on the empirical measurements, and (2) we need to upper bound a term of form $\sum_{k=1}^{K}\langle\bm{A}_{k},\bm{\rho}-\bm{\rho}^{\star}\rangle^{2}p_{k}$ originating from the exponential term in (3). We will address the first challenge by using a covering argument. For the second challenge, to utlize the upper bound for $\|\mathcal{A}(\bm{\rho}-\bm{\rho}^{\star})\|_{2}^{2}=\sum_{k=1}^{K}\langle\bm{A}_{k},\bm{\rho}-\bm{\rho}^{\star}\rangle^{2}$ in Section 2, we capture the largest probabilities by
@@
The proof is given in Appendix D. With exact or approximate spherical $t$-design POVMs ($t\geq 3$), Equation 18 ensures a stable recovery of the ground-truth state when the number of state copies $M=O(nd^{2}\overline{r}^{2}\log n/\epsilon^{2})$. Notably, this uniform guarantee is applied to all MPOs. It is worth noting that the $n$-qubit Clifford group comprises a unitary (not spherical) 3-design POVM [71, 72, 73]. Sampling Clifford circuits uniformly at random reproduces the first $3$ moments of the full unitary group endowed with the Haar measure. Unfortunately, unitary 3-designs are not precisely equivalent to spherical 3-designs, thus we cannot efficiently construct spherical 3-designs. However, this work [74] demonstrates one method to generate spherical 3-designs by extracting them from unitary designs. On the other hand, the SIC-POVM is not a $3$-design POVM, hence the assumption of "uniformly" distributed probabilities may be indispensable.
@@
$\|P_{\mathbb{S}_{+}}(\widehat{\bm{\rho}})-\bm{\rho}^{\star}\|_{F}$ $=$ $\|P_{\mathbb{S}_{+}}(\widehat{\bm{\rho}})-P_{\mathbb{S}_{+}}(\bm{\rho}^{\star})\|_{F}$ (19)
$\leq$ $\|\widehat{\bm{\rho}}-\bm{\rho}^{\star}\|_{F}\leq\epsilon,$
@@
$\|\widehat{\bm{\rho}}-\bm{\rho}^{\star}\|_{1}$ $\lesssim\frac{d\overline{r}\sqrt{(1+\delta)ns\gamma_{t}(\bm{\rho}^{\star})\log n}}{(1-\delta)\sqrt{M}}$ (20)
$+2\sqrt{s}\|\bm{N}\|_{F}+\|\bm{N}\|_{1}.$
@@
Compared to (21), the recovery upper bound in Theorem 2 is tight and optimal, up to logarithmic terms.
@@
4 Projected Gradient Descent with Guaranteed Convergence

### v4

# Sample-Efficient Quantum State Tomography for Structured Quantum States in One Dimension

@@ I Introduction
$\hskip-14.22636pt\bm{\rho}(i_{1}\cdots i_{n},j_{1}\cdots j_{n})=\bm{X}_{1}^{i_{1},j_{1}}\bm{X}_{2}^{i_{2},j_{2}}\cdots\bm{X}_{n}^{i_{n},j_{n}},$ (1)
@@ I-A Main contributions
TABLE I: Comparison with existing work for estimating $n$-qudits MPOs (with bond dimension $\overline{r}$) to achieve $\epsilon$-closeness in the Frobenius norm. The result for SIC/2-designs depends on $\gamma(\bm{\rho})=K\cdot\max_{k}p_{k}$, i.e., the probability distribution of measuring the state $\bm{\rho}$ with the corresponding POVM.
POVM #\# POVMs #\# total state copies
@@
SIC/22-designs (Theorem 1)        11 $O(\frac{nd^{2}\gamma(\bm{\rho})\overline{r}^{2}\log n}{\epsilon^{2}})$
$t$-designs, $t\geq 3$ (Theorem 2)        11 $O(\frac{nd^{2}\overline{r}^{2}\log n}{\epsilon^{2}})$
@@
We begin by studying recovery guarantees for MPOs from empirical quantum measurements (physical measurements containing statistical errors) obtained using a class of IC-PVOMs including the rank-one symmetric IC-POVM (SIC-POVM) [38] and more broadly spherical $t$-design POVMs that are induced by (approximate) spherical $t$-designs [39, 40]. The first contribution of this paper—presented in Section III—is a new sample complexity bound (refers to the number of state copies) that matches the number of parameters in MPOs to guarantee bounded recovery error for a particular estimator—the solution to a constrained least-squares optimization problem. Specifically, for SIC-POVM and $2$-design POVMs, our result shows that $M=O(nd^{2}\gamma(\bm{\rho})\overline{r}^{2}\log n/\epsilon^{2})$ number of state copies is sufficient to achieve $\epsilon$-accuracy in recovering the ground-truth state $\bm{\rho}$ in the Frobenius norm, where $\gamma(\bm{\rho})=K\cdot\max_{k=1}^{K}p_{k}$ captures the uniformity of the probabilistic distribution of the measurement. The term $\gamma(\bm{\rho})$ can be as small as 1, but in some cases, it may be as large as $d^{n}$; however, for many cases it is only polynomially large in $n$. The dependence of $\gamma(\bm{\rho})$ could be eliminated by using spherical $t$-designs with $t\geq 3$, for which we only need $M=O(nd^{2}\overline{r}^{2}\log n/\epsilon^{2})$ to ensure the sufficient condition. Excluding the logarithmic term, this bound exhibits linear dependence on $n$, $d^{2}$, and $\overline{r}^{2}$, aligning precisely with the degrees of freedom $O(nd^{2}\overline{r}^{2})$ in MPOs and is thus sample efficient. This result improves upon previous work [33] mentioned above using Haar random projective measurement bases that requires $\Omega(nd^{2}\overline{r}^{2}\log n)$ POVMs and $O(n^{3}d^{2}\overline{r}^{2}\log n/\epsilon^{2})$ total number of state copies. We summarize the main results and comparison with our previous result [33] in Table I. Moreover, physical MPO states are often approximately low-rank, allowing the previous analysis to be extended to the trace distance. Specifically, to ensure a bounded recovery error in terms of trace distance for rank-$s$ MPO states, the required number of state copies remains proportional to the degrees of freedom, $O(nsd^{2}\overline{r}^{2})$, when using (approximate) spherical $t$-designs with $t\geq 2$.
@@
The second contribution of this paper—presented in Section IV—is the development of a projected gradient descent (PGD) with guaranteed convergence analysis for solving the aforementioned constrained least-squares optimization problem. In each iteration, the algorithm performs a gradient update, followed by a projection back to the MPO format with a unit trace. Even without the constraint on the trace, computing the optimal projection onto the set of MPOs is challenging and could be NP-hard [41]. We employ a sequential approach called TT-SVD [27]—developed for computing TT decomposition which is equivalent to MPO—for computing an approximate projection. While TT-SVD is not optimal, it is computationally efficient and comes with a sub-optimal guarantee. Leveraging this guarantee, we show that, with an appropriate initialization, PGD converges linearly to a vicinity of the ground-truth state, with a distance matching the results in Table I. In addition, we show that spectral initialization provides a valid starting point, ensuring the convergence of PGD. Our numerical experiments in Section V suggest that PGD can also efficiently find the underlying state even with a random initialization.
@@
We note that our work is centered on the fundamental question of whether the number of state copies required for QST can be efficient (i.e. proportional to the intrinsic degrees of freedom) for recovering a structured quantum many-body state. We investigate this problem using IC-POVMs, focusing on sample complexity while not accounting for the implementation complexity of the measurement settings. In practice, the SIC-POVM or the spherical $t$-designs with $t\geq 2$ have no known efficient implementation using local quantum circuits. To partially address this practicality concern, we provide numerical evidence that local SIC-POVMs, which can be implemented efficiently and readily on current quantum hardware, can be used to perform QST for typical MPO states with $O(n)$ state copies using the proposed PGD algorithm in Section V.
I-B Notation
We use calligraphic letters (e.g., $\mathcal{X}$) to denote tensors, bold capital letters (e.g., $\bm{X}$) to denote matrices, bold lowercase letters (e.g., $\bm{x}$) to denote column vectors, and italic letters (e.g., $x$) to denote scalar quantities. Elements of matrices and tensors are denoted in parentheses. For example, $\mathcal{X}(i_{1},i_{2},i_{3})$ denotes the element in position $(i_{1},i_{2},i_{3})$ of the order-3 tensor $\mathcal{X}$. The calligraphic letter $\mathcal{A}$ is reserved for the linear measurement map. For a positive integer $K$, $[K]$ denotes the set $\{1,\dots,K\}$. The superscripts $(\cdot)^{\top}$ and $(\cdot)^{\dagger}$ denote the transpose and Hermitian transpose, respectively  [^3] As is conventional in the quantum physics literature (but not in information theory and signal processing), we use $(\cdot)^{\dagger}$ to denote the Hermitian transpose.. For two matrices $\bm{A},\bm{B}$ of the same size, $\left\langle\bm{A},\bm{B}\right\rangle=\operatorname{trace}(\bm{A}^{\dagger}\bm{B})$ denotes the inner product between them. $\|\bm{X}\|$ and $\|\bm{X}\|_{F}$ respectively represent the spectral norm and Frobenius norm of $\bm{X}$. $\sigma_{i}(\bm{X})$ is the $i$-th singular value of $\bm{X}$. For a vector $\bm{a}$ of size $N\times 1$, its $l_{n}$-norm is defined as $\bm{a}||_{n}=(\sum_{m=1}^{N}|a_{m}|^{n})^{\frac{1}{n}}$. For two positive quantities $a,b\in\mathbb{R}$, the inequality $b\lesssim a$ or $b=O(a)$ means $b\leq ca$ for some universal constant $c$; likewise, $b\gtrsim a$ or $b=\Omega(a)$ represents $b\geq ca$ for some universal constant $c$. To simplify notations in the following sections, for an order-$n$ TT format with ranks $(r_{1},\dots,r_{n})$, we define $\overline{r}=\max_{i=1}^{n-1}r_{i}$.
II Informationally Complete POVM
@@
$\hskip-17.07182pt\|\mathcal{A}(\bm{\rho})\|_{2}^{2}=\sum_{k=1}^{d^{2n}}\langle\bm{A}_{k},\bm{\rho}\rangle^{2}=\frac{\|\bm{\rho}\|_{F}^{2}+(\operatorname{trace}(\bm{\rho}))^{2}}{d^{n}(d^{n}+1)}.$ (3)
The proof is provided in Appendix A. By using tools from compressive sensing literature [57, 58, 59, 60, 61], the previous work [33] provides a similar stable embedding result from Haar random projective measurement bases, but only for MPOs. In contrast, with POVM being informationally complete, (3) holds for any Hermitian matrices without relying on any randomness in the measurement operator. This allows us to develop tight sample complexity bound in Section III and iterative algorithms with guaranteed convergence in Section IV. As an immediate consequence of Lemma 1, for any two density matrices $\bm{\rho}_{1}$ and $\bm{\rho}_{2}$, the difference $\bm{\rho}_{1}-\bm{\rho}_{2}$ remains Hermitian, and hence
@@
When $s=1$, we have $\frac{1}{K}\sum_{k=1}^{K}\bm{w}_{k}\bm{w}_{k}^{\dagger}=\int\bm{w}\bm{w}^{\dagger}d\bm{w}=\frac{1}{d^{n}}{\bf I}$, and hence $\bm{A}_{k}=\frac{d^{n}}{K}\bm{w}_{k}\bm{w}_{k}^{\dagger},k=1,\dots,K$ form a rank-one POVM. For simplicity, we call such an induced POVM $\{\bm{A}_{k}=\frac{d^{n}}{K}\bm{w}_{k}\bm{w}_{k}^{\dagger}\}$ as a $t$-design POVM. SIC-POVM is a special instance of $2$-design POVMs. In (5), we adopt the setting of uniform weights ($1/K$ for each $\bm{w}_{k}$, which is mostly commonly used in practice) to simplify the analysis, but one can consider a more general scenario of varied weights [40]. Unlike SIC-POVM, $t$-designs always exist and can, in principle, be constructed in any dimension and for any $t$ [64, 65], although in some cases these constructions can be inefficient, as they require vector sets of exponential size [66]. Instead, we may use approximate $t$-designs which can be efficiently implemented in practice [67].
@@
$(1-\delta)\int(\bm{w}\bm{w}^{\dagger})^{\otimes s}d\bm{w}\leq\frac{1}{K}\sum_{k=1}^{K}(\bm{w}_{k}\bm{w}_{k}^{\dagger})^{\otimes s}\leq(1+\delta)\int(\bm{w}\bm{w}^{\dagger})^{\otimes s}d\bm{w}$ (6)
@@
$(1-\delta)\frac{d^{n}(\|\bm{\rho}\|_{F}^{2}+(\operatorname{trace}({\bm{\rho}}))^{2})}{K(d^{n}+1)}\leq\|\mathcal{A}(\bm{\rho})\|_{2}^{2}\leq(1+\delta)\frac{d^{n}(\|\bm{\rho}\|_{F}^{2}+(\operatorname{trace}({\bm{\rho}}))^{2})}{K(d^{n}+1)}.$ (7)
The proof is given in Appendix B. Similar to (4), it follows from Equation 7 that for any two density matrices $\bm{\rho}_{1}$ and $\bm{\rho}_{2}$, $\mathcal{A}(\bm{\rho}_{1}-\bm{\rho}_{2})$ satisfies
@@
III Stable Recovery of Matrix Product Operators
@@
$\widehat{\bm{\rho}}=\operatorname*{\text{arg~min}}_{\bm{\rho}\in\mathbb{X}_{\bm{r}}}\|\mathcal{A}(\bm{\rho})-{\widehat{\bm{p}}}\|_{2}^{2},$ (10)
@@
$0$ $\!\!\!\!\leq\!\!\!\!$ $\|\mathcal{A}(\bm{\rho}^{\star})-\widehat{\bm{p}}\|_{2}^{2}-\|\mathcal{A}(\widehat{\bm{\rho}})-\widehat{\bm{p}}\|_{2}^{2}$ (12)
@@
$\operatorname{\mathbb{P}}\left(\sum_{k=1}^{K}a_{k}(\frac{f_{k}}{M}-p_{k})>t\right)\leq e^{-\frac{Mt}{4a_{\max}}\min\bigg\{1,\frac{a_{\max}t}{4\sum_{k=1}^{K}a_{k}^{2}p_{k}}\bigg\}}+e^{-\frac{Mt^{2}}{8\sum_{k=1}^{K}a_{k}^{2}p_{k}}},$ (15)
@@
Case $(i)$: 22-designs.
There are two potential technical difficulties in directly applying Equation 15 to bound the term $\langle\bm{\eta},\mathcal{A}(\widehat{\bm{\rho}}-\bm{\rho}^{\star})\rangle$ by plugging in $a_{k}=\langle\bm{A}_{k},\widehat{\bm{\rho}}-\bm{\rho}^{\star}\rangle$: (1) $a_{k}$ in not independent to the random variables $\{f_{k}\}$ since $\widehat{\bm{\rho}}$ depends on the empirical measurements, and (2) we need to upper bound a term of form $\sum_{k=1}^{K}\langle\bm{A}_{k},\bm{\rho}-\bm{\rho}^{\star}\rangle^{2}p_{k}$ originating from the exponential term in (15). We will address the first challenge by using a covering argument. For the second challenge, to utilize the upper bound for $\|\mathcal{A}(\bm{\rho}-\bm{\rho}^{\star})\|_{2}^{2}=\sum_{k=1}^{K}\langle\bm{A}_{k},\bm{\rho}-\bm{\rho}^{\star}\rangle^{2}$ in Section II, we capture the largest probabilities by
@@
The proof is given in Appendix D. With exact or approximate spherical $t$-design POVMs ($t\geq 3$), Equation 18 ensures a stable recovery of the ground-truth state when the number of state copies $M=O(nd^{2}\overline{r}^{2}\log n/\epsilon^{2})$. Notably, this uniform guarantee is applied to all MPOs. It is worth noting that the $n$-qubit Clifford group comprises a unitary (not spherical) 3-design POVM [71, 72, 73]. Sampling Clifford circuits uniformly at random reproduces the first $3$ moments of the full unitary group endowed with the Haar measure. Unfortunately, unitary 3-designs are not precisely equivalent to spherical 3-designs, thus we cannot efficiently construct spherical 3-designs. However, this work [74] demonstrates one method to generate spherical 3-designs by extracting them from unitary designs. On the other hand, the SIC-POVM is not a $3$-design POVM, hence the assumption of ”uniformly” distributed probabilities may be indispensable.
@@
$\|P_{\mathbb{S}_{+}}(\widehat{\bm{\rho}})-\bm{\rho}^{\star}\|_{F}=\|P_{\mathbb{S}_{+}}(\widehat{\bm{\rho}})-P_{\mathbb{S}_{+}}(\bm{\rho}^{\star})\|_{F}\leq\|\widehat{\bm{\rho}}-\bm{\rho}^{\star}\|_{F}\leq\epsilon,$ (19)
@@
$\|\widehat{\bm{\rho}}-\bm{\rho}^{\star}\|_{1}\lesssim\frac{d\overline{r}\sqrt{(1+\delta)ns\gamma_{t}(\bm{\rho}^{\star})\log n}}{(1-\delta)\sqrt{M}}+2\sqrt{s}\|\bm{N}\|_{F}+\|\bm{N}\|_{1}.$ (20)
@@
Compared to (21), the recovery upper bound in Theorem 2 is tight, up to logarithmic terms.
@@
IV Projected Gradient Descent with Guaranteed Convergence